Cash Payback
Anderson Company must evaluate two capital expenditure proposals. Anderson's hurdle rate is 12%. Data for the two proposals follow.

Proposal X Proposal Y
Required investment $300,000 $300,000
Annual after-tax cash inflows 30,000
After-tax cash inflows at the end of years 3, 6, 9, and 12 75,000
Life of project 12 years 12 years


What is the cash payback period for Proposal X? For Proposal Y?

Hint: For Proposal Y, in what year (3, 6, 9 or 12) will the full original investment be recovered?

Round Proposal X answer to one decimal place, if applicable.

Proposal X

Answer years

Proposal Y

Answer years
Answer #1

Payback Period = investment / Annula Cash inflows

Project X= 300,000 /30,000 = 10 Years

Project Y= 300000/75000 =4 Cash flows = 3,6, 9and 12 = 12Years

Answers :

Project X= 10 Years

Project Y = 12 Years